a rock is thrown horizontally from the top of a cliff 88m high, with a horizontal speed of 25m/s. with what velocity does the rock hit?


1
Expert's answer
2022-05-12T14:00:02-0400

"v_x=25~\frac ms,"

"v_y=\sqrt{2gh}=41.5~\frac ms,"

"v=\sqrt{v_x^2+v_y^2}=48.5~\frac ms."